OpenStudy (darkbluechocobo):

Check my answer please. Find the zeros of f(x) = x2- x - 6 by using a table and a graph. Enter the largest zero. Use a number only. I got -1 for the largest Zero

jimthompson5910 (jim_thompson5910):

Let's check that answer. Plug in x = -1 to get f(x) = x^2 - x - 6 f(-1) = (-1)^2 - (-1) - 6 ... replace EVERY x with -1 f(-1) = 1 - (-1) - 6 f(-1) = 1 + 1 - 6 f(-1) = 2 - 6 f(-1) = -4 Notice how the result is NOT zero, so that means that x = -1 is NOT a root or zero of f(x)

OpenStudy (darkbluechocobo):

Wouldn't they both be negative though (x-1) (x+6) so its 6 sorry

OpenStudy (darkbluechocobo):

wait no

jimthompson5910 (jim_thompson5910):

x^2 - x - 6 doesn't factor to (x-1) (x+6)

OpenStudy (darkbluechocobo):

Yes just realized.

OpenStudy (darkbluechocobo):

-3 and 2

jimthompson5910 (jim_thompson5910):

check each one

OpenStudy (darkbluechocobo):

3 -2

OpenStudy (darkbluechocobo):

because it works out that way

OpenStudy (darkbluechocobo):

-2^2= 4 4+2=6 6-6=0

OpenStudy (darkbluechocobo):

3^2=9 9-3=6 6-6=0

OpenStudy (darkbluechocobo):

so i believe the answer would be 3

jimthompson5910 (jim_thompson5910):

very good, you got it

OpenStudy (darkbluechocobo):

Thank you :D

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (darkbluechocobo):

Find the zeros of the function h(x) = 3x2 + 13x + 4 by factoring. 4, -1/3 4, 1/3 -4, 1/3 -4, -1/3 Ok I did some more but then I got to this one and was like Wahhhhhhhh

OpenStudy (darkbluechocobo):

Could you assist me

jimthompson5910 (jim_thompson5910):

if factoring is a pain, you could use the quadratic formula to solve